If the pOH of an aqueous NaOH solution is 11.55, what is the [H3O+] for this solution?


1
Expert's answer
2021-06-04T15:01:04-0400

Solution:

NaOH is a strong base, which means that it is 100% ionized in solution:

NaOH → Na+(aq) + OH(aq)

(pH must be greater than 7).


Therefore, most likely pH is 11.55, not pOH!!!


Assuming pH = 11.55, convert between pH and [H3O+] using the following equation:

pH = − log[H3O+]

[H3O+] = 10−pH

[H3O+] = 10−11.55 = 2.82×10−12

[H3O+] = 2.82×10−12 M


Answer: [H3O+] = 2.82×10−12 M